On a frictionless horizontal air table, puck A (with mass 0.248 kg ) is moving toward puck B (with mass 0.369 kg ), which is initially at rest. After the collision, puck A has velocity 0.123 m/s to the left, and puck B has velocity 0.647 m/s to the right.
Answer #1

apply the law of conservation of momentuma as

total Intial Momentum = total final momentum

i.e m1u1 + m2u2 = m1v1 + m2v2

here u2 = 0 and u1= ?

so

0.248 * u1 + 0 = -0.248 * 0.123 + (0.647 * 0.369)

u1 = 0.208 /0.248

u1= inital velcoity of mass1 = 0.84 m/s